OpenStudy (anonymous):

Name the property of real numbers illustrated by the equation 5+{15+6}={5+15}+6 answers are associative property addition. Communicative property multiplication. Distributive property. Associative property multiplication

jigglypuff314 (jigglypuff314):

associative is when the parenthesis get moved around a+(b+c) = (a+b)+c (if multiplication) a*(b*c) = (a*b)*c communicative is when the terms get moved around a+b+c = b+a+c Distributive property a(b+c) = ab + ac